# Reference-Free Sampling-Based Model Predictive Control

### v1

@@ III-D Best trajectory tracking
One distinction is the separation between trajectory evolution and action execution. While the nominal trajectory $\tau_{0}$ evolves through standard MPPI importance-weighted averaging across iterations, the robot always executes actions from $\tau_{best}$, which is the best-performing trajectory that has been explicitly tested through simulation rollout. This serves two purposes: (1) it ensures that executed actions always come from a verified, fully-simulated trajectory rather than from an untested weighted mixture, providing safety guarantees; and (2) it prevents performance degradation during iterative refinement by maintaining monotonic improvement in executed solution quality.
@@ III-E Real-time state prediction and warm-starting strategy
Our approach differs from prior works that handle computation delays by constraining the first $\lfloor\Delta t/dt\rfloor$ actions across all parallel rollouts to match the actions being executed during computation. While conceptually simpler, this wastes computational budget by forcing identical initial actions across all sampled trajectories, reducing exploration diversity. Instead, our state prediction strategy optimizes from the predicted future state, allowing all samples to explore freely from the start of their planning horizons, thereby maximizing the effective use of the limited sample budget.
@@ IV-A Experimental setup
Experiments in simulation (Go2 quadruped and G1 humanoid). To complement hardware experiments, we use the Simple physics simulator [carpentier2024compliantrigidcontactsimulation], which provides high-fidelity, accurate, and efficient frictional contact dynamics simulation by proper handling of nonlinear complementarity contact models [brogliato:hal-01349852]. Simulation experiments serve two purposes: (i) showing behaviors that are unsafe or impractical to test on hardware, such as backflips, aggressive jumps, and high-speed locomotion up to $2.0\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$, and (ii) demonstrating cross-platform generalization by validating our method on the Unitree G1 humanoid.
@@ IV-B Motion discovery
A central contribution of our work is demonstrating reference-free motion discovery in real-time across different tasks and even robot morphologies. We systematically evaluate the quadruped’s ability to discover motion strategies purely from high-level goals and demonstrate that our framework transfers to a humanoid robot, enabling the achievement of walking gaits without requiring the tuning of gait parameters or other motion priors.
@@ IV-B1 Velocity-adaptive gaits
In simulation and on the real Go2 robot, we specify the desired forward velocity in the cost function without providing any explicit gait references or contact sequence. We use an extended planning horizon of $0.9\text{\,}\mathrm{s}$ that is important for locomotion discovery: contrary to reference-tracking methods that can use shorter horizons of $0.4\text{\,}\mathrm{s}$ following predefined patterns, emergent gait discovery requires sufficient horizon time to evaluate locomotion stability and quality. For slow walking in particular, a $0.9\text{\,}\mathrm{s}$ window allows the optimization to assess whether a motion pattern leads to stable periodic locomotion or results in falling. We evaluate three representative velocity commands:
@@
To analyze the different gaits, we plot the obtained contact patterns in Fig. 5(b) where the initial target velocity of $0.5\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ was increased linearly until reaching $2.0\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$.
@@ IV-B3 Jumping
We demonstrate different jumping behaviors by specifying only sparse task-level objectives where the goal is to jump vertically, turn the base, and land safely.
@@
As soon as the terminal height objective becomes visible in the current MPC horizon (see orange dashed line in Fig. 6), the robot initiates a multi-phase jumping maneuver. First, it crouches down by flexing its legs to then rapidly extends its legs to launch into the air, achieving a peak height of $0.7\text{\,}\mathrm{m}$ above the ground. During flight, the robot controls its body orientation to maintain stability and prepare for landing. We have successfully validated this jumping behavior on the real Go2 robot platform, as shown in Fig. 1(b).
Backflip: By specifying a target pitch orientation ($180\text{\,}\mathrm{\SIUnitSymbolDegree}$) at the highest point, the robot discovers a complete backflip maneuver. It coordinates a pre-jump crouch, explosive takeoff generating both vertical impulse and angular momentum, mid-air rotation control, and precise landing absorption to achieve the desired orientation and height. Starting from the elevated platform of $0.5\text{\,}\mathrm{m}$, the robot completes a full 360° rotation with landing as shown in Fig. 8(b).
IV-B4 Dynamic handstand balancing
Starting from a standard quadrupedal stance with a level base, when commanded to achieve an inverted pose ($45\text{\,}\mathrm{\SIUnitSymbolDegree}$ pitch angle), the robot executes a dynamic swing maneuver to transition into a handstand configuration. Then, the controller discovers stabilizing strategies through continuous leg repositioning and contact pattern adjustments. Fig. 7 illustrates simulation snapshots of emergent contact-making and breaking patterns, where the legs dynamically reposition to maintain balance. The accompanying plot shows that the handstand pose is stably maintained for about $1\text{\,}\mathrm{s}$ with low orientation error.
@@ IV-B5 Humanoid locomotion
To highlight the generality of our reference-free framework beyond quadrupedal locomotion, we also evaluate it on the G1 humanoid in simulation. Despite the shift in morphology and increase in DoFs (from 12 to 37), the same algorithm successfully discovers a walking gait from the exact high-level cost without modification, as displayed with overlayed snapshots in Fig. 8(a). While all previously demonstrated behaviors were achieved in real-time, the humanoid experiment requires an increase from 30 to 70 parallel sample trajectories due to the increased complexity of bipedal balance control, which prevents real-time execution. This cross-platform study underlines a key advantage of our approach: the identical framework, cost functions, and spline parameterization transfer across robot morphologies, enabling rapid deployment on new platforms.
@@ IV-C Ablation studies
(1) Dual-space Hermite spline parameterization. We compare our cubic Hermite parameterization with velocity targets against simpler alternatives: cubic splines without velocity targets and quadratic splines. The goal is to quantify the improvement in trajectory quality and stability that results from explicitly controlling both position and velocity at spline nodes. Note that we are using the best trajectory tracking mechanism III-D in all cases.
(2) Best trajectory tracking. We study the effect of executing actions from the best sampled trajectory $\tau_{best}$ rather than from the evolving nominal trajectory $\tau_{0}$. This ablation tests whether explicitly safeguarding execution against untested mixtures yields more reliable performance. In both cases, trajectories are parameterized using Hermite splines. We then track how often the optimizer failed to find an improvement over the shifted previous solution: walking $(27.9\%)$, jumping $(29.7\%)$, and handstand $(27.1\%)$.
For both ablations, we evaluate performance across multiple tasks, walking, handstands, and jumping, using ten randomized seeds per setting. The primary metric is the average trajectory cost, which directly reflects the optimization objective. In Tab. III, we report mean $\pm$ standard deviation to highlight consistency across runs. Failures report when the robot base hits the ground. For the handstand, we also plot the base rotation trajectories. For the jumping task, we report the mean $\pm$ standard deviation of the maximum reached height to provide more interpretable performance metrics. Across all tasks, our Hermite spline parameterization combined with best-trajectory tracking yields lower mean trajectory costs and less variance. This advantage is particularly evident in jumping, where our method achieves the highest mean height, and in handstands, where cubic and quadratic splines exhibit larger variance and occasional failures.
TABLE III: Ablation study table with figure. The success rate (out of 10 seeds) and mean $\pm$ standard deviation cost are reported for Walking and Handstand tasks. For the Jumping task, the mean $\pm$ standard deviation of maximum reached height is reported. Right. Comparison of base rotation trajectories during the Handstand task.
@@
Cubic (no vel) 6/10 1124.1 $\pm$ 135.7 10/10 1540.2 $\pm$ 655.8 63.8 $\pm$ 0.4
Quadratic (no vel) 0/10 9/10 1882.2 $\pm$ 493.9 47.7 $\pm$ 0.4
Best-trajectory (ours) 10/10 913.9 $\pm$ 25.8 10/10 811.7 $\pm$ 253.4 68.4 $\pm$ 0.2
Nominal-only 10/10 924.9 $\pm$ 24.3 10/10 844.1 $\pm$ 203.2 67.4 $\pm$ 0.3
@@ V Discussion and Conclusion
We presented a sampling-based MPC framework that enables reference-free locomotion by combining dual-space spline parameterization with diffusion-inspired noise annealing. In contrast to prior MPPI-inspired approaches that rely on thousands of GPU rollouts and/or gait priors, our method achieves real-time performance on standard CPU hardware with as few as 30 samples. Experiments on the Go2 quadruped demonstrate real-world behaviors ranging from adaptive gaits to jumping, while simulation studies showcase more complex acrobatics such as 180-degree turns, backflips, and handstand balancing. Additionally, successful transfer to the G1 humanoid in simulation demonstrates the discovery of bipedal walking without modification. These results highlight that diverse and dynamic locomotion strategies can emerge purely from high-level objectives, without reference tracking or pre-computed contact sequences.
By reducing computational demands and removing the need for handcrafted priors, our work demonstrates that a wide variety of locomotion strategies can emerge in real time. At the same time, the motions we obtain are not always fully optimized in terms of smoothness or efficiency. This limitation is an expected consequence of our design choice. Sampling spline control points in joint space and delegating torque-level execution to a PD controller reduces the search space, but also limits fine-grained optimization, since we do not optimize direct torque commands at every control step.
Rather than a drawback, we view this as a tradeoff that makes the approach particularly useful as a motion discovery mechanism. The framework provides a flexible source of candidate behaviors, contact sequences, and strategies that can serve as warm starts for higher-level solvers or be integrated into larger control pipelines. Future directions include improving the noise sampling scheme and combining our method with trajectory optimization or learning-based refinement to transform diverse exploratory motions into polished, task-specific controllers.
@@ Acknowledgments
This work has received support from the French government, managed by the National Research Agency, under the France 2030 program with the references Organic Robotics Program (PEPR O2R) and “PR[AI]RIE-PSAI” (ANR-23-IACL-0008). This research was funded, in part, by l’Agence Nationale de la Recherche (ANR), projects RODEO (ANR-24-CE23-5886) and PEPR O2R – PI3 ASSISTMOV (ANR-22-EXOD-0004). The European Union also supported this work through the ARTIFACT project (GA no.101165695) and the AGIMUS project (GA no.101070165). The Paris Île-de-France Région also supported this work in the frame of the DIM AI4IDF. Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the funding agencies.

### v2

@@ III-D Best trajectory tracking
One distinction is the separation between trajectory evolution and action execution. While the nominal trajectory $\tau_{0}$ evolves via standard MPPI importance-weighted averaging across iterations, the robot always executes actions from $\tau_{best}$, the best-performing trajectory tested in simulation rollout. This serves two purposes. First, it ensures that executed actions always come from a verified, fully-simulated trajectory rather than from an untested weighted mixture, providing safety guarantees. Second, it prevents performance degradation during iterative refinement by maintaining monotonic improvement in the quality of the executed solution.
@@ III-E Real-time state prediction and warm-starting strategy
Our approach differs from prior work that handles computation delays by constraining the first $\lfloor\Delta t/dt\rfloor$ actions across all parallel rollouts to match the actions executed during computation. While conceptually simpler, this wastes computational budget by forcing identical initial actions across all sampled trajectories, reducing exploration diversity. Instead, our state prediction strategy optimizes from the predicted future state, allowing all samples to explore freely from the start of their planning horizons, thereby maximizing the effective use of the limited sample budget.
@@ IV-A Experimental setup
Experiments in simulation (Go2 quadruped and G1 humanoid). To complement hardware experiments, we use the Simple physics simulator [5], which provides accurate and efficient frictional contact dynamics simulation by proper handling of nonlinear complementarity contact models [3]. Simulation experiments serve two purposes: (i) showing behaviors that are unsafe or impractical to test on hardware, such as backflips, aggressive jumps, and high-speed locomotion up to $2.0\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$, and (ii) demonstrating cross-platform generalization by validating our method on the Unitree G1 humanoid.
@@ IV-B Motion discovery
A central contribution of our work is the demonstration of reference-free motion discovery in real time across diverse tasks and robot morphologies. We evaluate the quadruped’s ability to discover motion strategies purely from high-level goals and show that our framework generalizes to a humanoid robot, enabling walking without tuning gait parameters or other motion priors. For the more aggressive behaviors shown in simulation, we do not claim fully optimized or hardware-ready motions, but rather proof-of-concept behaviors obtained from the same generic framework.
@@ IV-B1 Velocity-adaptive gaits
In simulation and on the real Go2 robot, we specify the desired forward velocity in the cost function without providing any explicit gait references or contact sequence. We use an extended planning horizon of $0.9\text{\,}\mathrm{s}$ that is important for locomotion discovery. In contrast to reference-tracking methods that can use shorter horizons of $0.4\text{\,}\mathrm{s}$ to follow predefined patterns, emergent gait discovery requires sufficient horizon time to evaluate locomotion stability and quality. For slow walking in particular, a $0.9\text{\,}\mathrm{s}$ window allows the optimization to assess whether a motion pattern leads to stable periodic locomotion or to a fall. We evaluate three representative velocity commands:
@@
To analyze the different gaits, Fig. 5(a) shows base-velocity tracking as the commanded speed is ramped from $0.5\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ to $2.0\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$. Fig. 5(b) shows the contact patterns, highlighting the transition from trotting to more dynamic gaits as the velocity command increases. The resulting behavior on the real Go2 platform is illustrated in Fig. 2.
@@ IV-B3 Jumping
We demonstrate diverse jumping behaviors by specifying only sparse task-level objectives where the goal is to jump vertically, turn the base, and land safely.
@@
As soon as the terminal height objective becomes visible in the current MPC horizon (see orange dashed line in Fig. 6), the robot initiates a multi-phase jumping maneuver. First, it crouches by flexing its legs, then rapidly extends them to launch into the air, reaching a peak height of $0.7\text{\,}\mathrm{m}$ above the ground. During flight, the robot controls its body orientation to maintain stability and prepare for landing. We have successfully validated this jumping behavior on the real Go2 robot platform, as shown in Fig. 1(b).
Backflip: By specifying a target pitch orientation ($180\text{\,}\mathrm{\SIUnitSymbolDegree}$) at the highest point, the robot discovers a complete backflip maneuver. It coordinates a pre-jump crouch, an explosive takeoff that generates both vertical impulse and angular momentum, mid-air rotation control, and precise landing absorption to achieve the desired orientation and height. Starting from the elevated platform of $0.5\text{\,}\mathrm{m}$, the robot completes a full 360° rotation with landing as shown in Fig. 8(b).
@@
IV-B4 Dynamic handstand balancing
Starting from a standard quadrupedal stance with a level base, when commanded to achieve an inverted pose ($45\text{\,}\mathrm{\SIUnitSymbolDegree}$ pitch angle), the robot executes a dynamic swing maneuver to transition into a handstand configuration. Then, the controller identifies stabilizing strategies by continuously repositioning the legs and adjusting contact patterns. Fig. 7 illustrates simulation snapshots of emergent contact-making and breaking patterns, where the legs dynamically reposition to maintain balance. The accompanying plot shows that the handstand pose is stably maintained for about $1\text{\,}\mathrm{s}$ with low orientation error.
@@ IV-B5 Humanoid locomotion
To highlight the generality of our reference-free framework beyond quadrupedal locomotion, we also evaluate it on the G1 humanoid in simulation. Despite the shift in morphology and increase in DoFs (from 12 to 37), the same algorithm successfully discovers a walking gait from the exact high-level cost without modification, as displayed with overlayed snapshots in Fig. 8(a). While all previously demonstrated behaviors were achieved in real-time, the humanoid experiment requires an increase from 30 to 70 parallel sample trajectories due to the increased complexity of bipedal balance control, which prevents real-time execution. This cross-platform study underlines a key advantage of our approach. The identical framework, cost functions, and spline parameterization transfer across robot morphologies, enabling rapid deployment on new platforms. However, these humanoid results are currently limited to simulation and should be interpreted as a proof of concept for motion discovery rather than as hardware-ready whole-body behaviors. The discovered motions are not yet optimized for smoothness, efficiency, or robustness to real-world sensing and actuation constraints.
@@ IV-C Ablation studies
(1) Cubic Hermite splines. We compare our cubic Hermite spline parameterization with velocity targets against cubic splines without velocity targets and quadratic splines. The goal is to quantify the improvement in trajectory quality and stability achieved by explicitly controlling both position and velocity at spline nodes. Note that we are using the best trajectory tracking mechanism III-D in all cases.
(2) Best trajectory tracking. We study the effect of executing actions from the best sampled trajectory $\tau_{best}$ rather than from the evolving nominal trajectory $\tau_{0}$. This ablation tests whether explicitly safeguarding execution against untested mixtures yields more reliable performance. In both cases, trajectories are parameterized using Hermite splines. We then track how often the optimizer failed to improve upon the shifted previous solution: walking $(27.9\%)$, jumping $(29.7\%)$, and handstand $(27.1\%)$.
For both ablations, we evaluate performance across multiple tasks, walking, handstand, and jumping, using 10 randomized seeds per setting. The primary metric is the average trajectory cost, which directly reflects the optimization objective. In Tab. III, we report mean $\pm$ standard deviation to highlight consistency across runs. Failures are recorded when the robot base hits the ground. For the handstand task, Fig. 9 additionally shows the base pitch trajectories. For the jumping task, we report the mean $\pm$ standard deviation of the maximum reached height to provide a more interpretable performance metric. Across all tasks, our Hermite spline parameterization, combined with best-trajectory tracking, yields lower mean trajectory costs and lower variance. This advantage is particularly evident in jumping, where our method achieves the highest mean height (h), and in handstands, where cubic and quadratic splines exhibit larger variance and occasional failures.
TABLE III: Ablation results. Values are success rates (over 10 seeds) and mean $\pm$ std for continuous metrics.
@@
Cubic 6/10 1124.1 $\pm$ 135.7 10/10 1540.2 $\pm$ 655.8 63.8 $\pm$ 0.4
Quadratic 0/10   – 9/10 1882.2 $\pm$ 493.9 47.7 $\pm$ 0.4
Best traj. (ours) 10/10 913.9 $\pm$ 25.8 10/10   811.7 $\pm$ 253.4 68.4 $\pm$ 0.2
Nominal only 10/10 924.9 $\pm$ 24.3 10/10   844.1 $\pm$ 203.2 67.4 $\pm$ 0.3
Figure 9: Handstand ablation: base pitch trajectories for different spline parameterizations during the handstand task.
@@ V Discussion and Conclusion
We presented a sampling-based MPC framework that enables reference-free locomotion by combining cubic Hermite spline parameterization with diffusion-inspired noise annealing. In contrast to prior MPPI approaches that rely on thousands of GPU rollouts and gait priors, our method achieves real-time performance on a CPU with as few as 30 samples. Experiments on the Go2 quadruped and G1 humanoid show that the same framework can produce diverse locomotion behaviors across tasks and robot morphologies without reference tracking or predefined contact sequences.
A limitation of the current formulation is that the resulting motions are not always fully optimized with respect to smoothness or efficiency. This follows from our design choice to sample spline control points in joint space and delegate torque-level execution to a PD controller, which reduces the search space but also limits fine-grained optimization compared with direct torque optimization.
This is a drawback of the present formulation, but also a trade-off that makes the approach useful as a motion-discovery mechanism. The framework provides a flexible source of candidate behaviors, contact sequences, and strategies that can serve as warm starts for higher-level solvers or be integrated into larger control pipelines. Future directions include improving the noise sampling scheme and integrating our method with trajectory optimization or learning-based refinement to transform diverse exploratory motions into polished, task-specific controllers.
@@ Acknowledgments
Supported by the French government via ANR under France 2030 (PEPR O2R, PR[AI]RIE-PSAI ANR-23-IACL-0008, RODEO ANR-24-CE23-5886); the EU (ARTIFACT 101165695, AGIMUS 101070165); and Région Île-de-France (DIM AI4IDF). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the funding agencies.